# Smooth orbit equivalence rigidity for dissipative geodesic flows

### v1

@@ 3. Action on holomorphic differentials
We have seen that, by passing through a specific type of $2$-currents instead of directly using the pullback $\phi^{*}$, the linear map $\Phi:\mathcal{D}^{\prime}_{\text{tr}}(SM)\to\mathcal{D}^{\prime}_{\text{tr}}(\tilde{S}M)$ defined in (2.3) sends distributional solutions to the transport equation of one generalized thermostat to those of the second. In this section, we want to show that $\Phi$ can also be seen as acting on holomorphic differentials from one complex surface to another when $\lambda$ is of Fourier degree $\leq 1$ and the attenuated tensor tomography problem of order $1$ is satisfied.
3.1. Action on fiberwise holomorphic distributions
We start by studying the action of $\Phi$ on the subspace $\mathcal{D}^{\prime}_{\text{tr},+}(SM)$ defined in (2.6). This will require some microlocal analysis.
We introduce $\mathcal{C}\subseteq\{(y,\eta)\in T^{*}(SM)\mid\eta(F(y))=0\}$, the closed cone enclosed by $(E^{s})^{\ast}$ and $(E^{u})^{\ast}$ in the half-space $\{(y,\eta)\in T^{*}(SM)\mid\eta(V(y))\geq 0\}$. See Figure 1.
@@ 3.1. Action on fiberwise holomorphic distributions
Let $(M,g,\lambda)$ be an Anosov generalized thermostat. If $u\in\mathcal{D}^{\prime}_{\textup{tr},+}(SM)$, then $\textup{WF}(u)\subseteq\mathcal{C}$ and $u_{k}\in\Omega_{k}$ for all $k\in\mathbb{Z}$.
@@
The argument is essentially the same as that of [GLP24, Lemma 2.5]. Let us give the details.
@@
$\text{WF}(u)=\text{WF}(S_{+}u)\subset\{(y,\eta)\in T^{*}(SM)\mid\eta(V(y))\geq 0\}.$
Given that $(F+V(\lambda))u=0$, elliptic regularity tells us that
$\text{WF}(u)\subset\Sigma=\{(y,\eta)\in T^{*}(SM)\mid\eta(F(y))=0\}.$
By propagation of singularities for real principal type differential operators (see [Hö09, Theorem 26.1.1]), we further know that $\text{WF}(u)$ is invariant by the symplectic lift of the flow $\varphi_{t}$. Given the Anosov property, the maximal flow-invariant subset of $T^{*}(SM)$ contained in $\{\eta(F(y))=0,\,\eta(V(y))\geq 0\}$ is $\mathcal{C}$, so this gives us the first claim.
For the second claim, recall that $u_{k}=(2\pi)^{-1}\pi_{k}^{\ast}\pi_{k\ast}u$. The pushforward operator $\pi_{k\ast}$ only selects the wavefront set in $(\mathbb{R}F)^{\ast}\oplus\mathbb{H}^{*}$ (see [FT99, Proposition 11.3.3]), which is empty given that $\mathcal{C}\cap((\mathbb{R}F)^{\ast}\oplus\mathbb{H}^{*})=\{0\}$ by property (2.11). Therefore $u_{k}\in\Omega_{k}$.
We will also need the following lemma with the same proof as [GLP24, Lemma 3.3].
@@
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$ be Anosov generalized thermostats. Suppose there exists a smooth orbit equivalence $\phi:\tilde{S}M\to SM$ isotopic to the identity between them. Then $\phi$ preserves the natural orientation of the weak unstable bundle, namely that given by the basis $\{F,Y^{u}\}$.
Armed with this, we can show that $\Phi$ maps fiberwise holomorphic distributional solutions to the transport equation of one thermostat to those of the second. For this step of the proof, however, we restrict to thermostats where $\lambda$ has Fourier degree $\leq 1$ and the attenuated tensor tomography problem of order $1$ is satisfied.
@@
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$, with $\lambda$ and $\tilde{\lambda}$ of Fourier degree $\leq 1$, be Anosov thermostats satisfying the attenuated tensor tomography problem of order $1$. Suppose there exists a smooth orbit equivalence $\phi:\tilde{S}M\to SM$ isotopic to the identity between them. Then the map $\Phi$ defined in (2.3) yields a $\mathbb{C}$-linear isomorphism
$\Phi:\mathcal{D}^{\prime}_{\textup{tr},+}(SM)\to\mathcal{D}^{\prime}_{\textup{tr},+}(\tilde{S}M).$
@@
Since $d\phi^{T}$ maps connected sets to connected sets, $(E^{u})^{*}$ to $(\tilde{E}^{u})^{*}$, and $(E^{s})^{*}$ to $(\tilde{E}^{s})^{*}$, $d\phi^{T}(\mathcal{C})$ must be one of the four cones depicted on the right of Figure 1 (inside the characteristic set $\tilde{\Sigma}$). It follows that $d\phi^{T}(\mathcal{C})=\pm\tilde{\mathcal{C}}$ because any other cone would entail that $\phi$ reverses the orientation, which is impossible since it is assumed to be isotopic to the identity. If $d\phi^{T}(\mathcal{C})=-\tilde{\mathcal{C}}$, then $\phi$ would flip the orientation of the weak unstable leaves, which contradicts Lemma 3.2. Therefore $d\phi^{T}(\mathcal{C})=\tilde{\mathcal{C}}$.
Let $u\in\mathcal{D}^{\prime}_{\textup{tr},+}(SM)$ and $\tilde{u}:=\Phi u$. By Lemma 3.1, we know that $\text{WF}(u)\subseteq\mathcal{C}$. By Lemma 2.4, we thus know that $\text{WF}(\tilde{u})=d\phi^{T}(\text{WF}(u))\subseteq\tilde{\mathcal{C}}$. Then $\tilde{S}_{-}\tilde{u}\in\mathcal{C}^{\infty}(\tilde{S}M)$ and, since $(\tilde{F}+\tilde{V}(\tilde{\lambda}))\tilde{u}=0$, we also have that
$(\tilde{F}+\tilde{V}(\tilde{\lambda}))\tilde{S}_{-}\tilde{u}$ $=[\tilde{F}+\tilde{V}(\tilde{\lambda}),\tilde{S}_{-}]\tilde{u}$
$=\left(\tilde{\eta}_{+}+\tilde{\lambda}_{1}\tilde{V}+i\tilde{\lambda}_{1}\right)\tilde{u}_{0}-\left(\tilde{\eta}_{-}+\tilde{\lambda}_{-1}\tilde{V}-i\tilde{\lambda}_{-1}\right)\tilde{u}_{1}.$
Since $(M,\tilde{g},\tilde{\lambda})$ satisfies the tensor tomography problem of order $1$ by assumption, it follows that $\tilde{S}_{-}\tilde{u}$ is of Fourier degree 00. Hence $\tilde{u}$ is fiberwise holomorphic. The fact that $\Phi$ is an isomorphism is then clear as it admits an inverse, namely the map associated to $(\phi^{-1})^{*}$ by (2.3). ∎
@@ 3.2. Extension operator
is well-defined. Indeed, if $u\in\mathcal{D}^{\prime}_{\text{tr},+}(SM)$, then $Xu+V(\lambda u)=(F+V(\lambda))u=0$, which means that $(Xu)_{0}=0$ and hence $\eta_{-}u_{1}=0$. By Lemma 2.5, this is equivalent to ¯$\overline{\partial}\pi_{1\ast}u=0$.
@@
$\Psi:H^{0}_{J}(M,\kappa)\to H^{0}_{\tilde{J}}(M,\kappa)$
by the commutative diagram
${\mathcal{D}^{\prime}_{\text{tr},+}(SM)}$${\mathcal{D}^{\prime}_{\text{tr},+}(\tilde{S}M)}$${H^{0}_{J}(M,\kappa)}$${H^{0}_{\tilde{J}}(M,\kappa).}$$\scriptstyle{\Phi}$$\scriptstyle{\pi_{1\ast}}$$\scriptstyle{e_{1}}$$\scriptstyle{\Psi}$ (3.2)
@@ 3.3. Period preservation
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$, with $\lambda$ and $\tilde{\lambda}$ of Fourier degree $\leq 1$, be Anosov thermostats satisfying the attenuated tensor tomography problem of order $1$. The $\mathbb{C}$-linear map
$\Psi:H^{0}_{J}(M,\kappa)\to H^{0}_{\tilde{J}}(M,\kappa)$
@@
Recall that there is a push-forward map $\pi_{\ast}:\mathcal{C}^{\infty}(SM,\Omega^{2}(SM))\to\mathcal{C}^{\infty}(M,\Omega^{1}(M))$ given by integration along fibers. It satisfies $d\pi_{\ast}=\pi_{\ast}d$ (see [BT82, Proposition 6.14]), and it extends to currents. By [BT82, Proposition 6.15], we have the projection formula
@@
Let $(M,g,\lambda)$ be a generalized thermostat. For any $u\in\mathcal{D}^{\prime}(SM)$, we have
@@
It suffices to establish the claim for $u\in\mathcal{C}^{\infty}(SM)$. Recall that $\omega:=\iota_{F}\mu$. A quick computation then yields
@@
$\pi_{\ast}(u\omega)_{x}(w)=\int_{S_{x}M}\iota_{\widetilde{w}}(u\omega),$
where $\widetilde{w}\in T(SM)$ is a lift of $w$ under $d\pi$. We take $\widetilde{w}=(w,0)$, i.e., no component in the subbundle $\mathbb{V}=\ker d\pi$. Then, since $\psi(\tilde{w})=0$, we get
$\pi_{\ast}(u\omega)_{x}(w)=\int_{S_{x}M}\iota_{\widetilde{w}}(\beta)u\psi.$
@@
where in the last equality we used the formula (2.4). In terms of $1$-forms, since $u_{k}=(2\pi)^{-1}\pi_{k}^{*}\pi_{k\ast}u$, we proved that
@@
We conclude by applying $\star$ to both sides.
@@
Let $(M,g,\lambda)$ be an Anosov generalized thermostat and $\gamma$ a closed thermostat geodesic. For any $u\in\mathcal{D}^{\prime}_{\textup{tr}}(SM)$, the pairing $\langle\pi^{-1}(\gamma),u\omega\rangle$ is well-defined and
@@
By the wavefront set calculus, the pairing $\langle\pi^{-1}(\gamma),u\omega\rangle$ is well-defined whenever
@@
(see [Hö03, Corollary 8.2.7] for instance). The conormal bundle $N^{*}(\pi^{-1}(\gamma))$ consists of a line contained in $(\mathbb{R}F)^{*}\oplus\mathbb{H}^{*}$, so Lemma 3.1 and property (2.11) tell us that the intersection with $\text{WF}(u)$ is indeed empty. It follows that the pairing $\langle\pi^{-1}(\gamma),u\omega\rangle$ is well-defined and extends the pairing computed for $u\in\mathcal{C}^{\infty}(SM)$.
We can then apply the projection formula (3.3) and Lemma 3.5. As seen in Lemma 2.3, the $2$-current $u\omega$ is closed if $u\in\mathcal{D}^{\prime}_{\text{tr}}(SM)$, so $\star(\pi_{-1\ast}u+\pi_{1\ast}u)$ is also closed given that $\pi_{*}d=d\pi_{*}$, which implies that its integral only depends on the homology class $[\gamma]$.
As $\pi_{1}^{*}\star=-V\pi_{1}^{*}$, for any $u\in\mathcal{D}^{\prime}(SM)$ we may write
@@
Therefore, if $\tau\in H^{0}_{J}(M,\kappa)$, $[\gamma]\in H_{1}(M;\mathbb{Z})$, and $\gamma$ is any thermostat geodesic whose homology class is $[\gamma]$, Lemma 3.6 gives us
@@
Let $[\gamma]\in H_{1}(M;\mathbb{Z})$ and let $\gamma$, $\tilde{\gamma}$ be two thermostat geodesics (with respect to $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$) whose homology class is $[\gamma]$. Since $\phi$ is isotopic to the identity, we know that $[\phi(\pi^{-1}(\tilde{\gamma}))]=[\pi^{-1}(\gamma)]$ in $H_{2}(SM;\mathbb{Z})$.
We claim that the pairing $\langle\phi(\pi^{-1}(\tilde{\gamma})),e_{1}(\tau)\omega\rangle$ is well-defined. The tangent space to $\pi^{-1}(\tilde{\gamma})$ is $(\mathbb{R}\tilde{F})\oplus\mathbb{\tilde{V}}$. By property (2.9), it trivially intersects the closed cone ~$\tilde{\mathcal{B}}\subseteq\tilde{E}^{s}\oplus\tilde{E}^{u}$ enclosed by $\tilde{E}^{s}$ and $\tilde{E}^{u}$, and whose orthogonal projection onto ~$\tilde{\mathbb{H}}\oplus\tilde{\mathbb{V}}$ avoids $\tilde{\mathbb{V}}$. Since $\phi$ preserves the orientation, the same arguments as in Proposition 3.3 tell us that the tangent space to $\phi(\pi^{-1}(\tilde{\gamma}))$ intersects the closed cone $\mathcal{B}$ trivially. As a result, its conormal $N^{*}(\phi(\pi^{-1}(\tilde{\gamma})))$ avoids the closed cone $\mathcal{C}$, which contains $\text{WF}(e_{1}(\tau))$ by Lemma 3.1. The wavefront set condition (3.5) is hence satisfied.
The $2$-current $e_{1}(\tau)\omega$ is closed by Lemma 2.3. By the Hodge decomposition theorem, we may hence write $e_{1}(\tau)\omega=\sigma+df$ for some harmonic $2$-current $\sigma$ and $1$-current $f$ with $\text{WF}(f)=\text{WF}(e_{1}(\tau))$. Thanks to the wavefront set condition, the same argument as for $e_{1}(\tau)\omega$ then shows that both pairings $\langle\pi^{-1}(\gamma),df\rangle$ and $\langle\phi(\pi^{-1}(\tilde{\gamma})),df\rangle$ are well-defined. They must be equal to 00 since $df$ is exact. We thus get
@@
where in the second equality we have used the fact that $\sigma$ is harmonic and $[\pi^{-1}(\gamma)]=[\phi(\pi^{-1}(\tilde{\gamma}))]$ in $H_{2}(SM;\mathbb{Z})$.
We can now use (3.6) and unravel the definitions to obtain
@@
$=2i\int_{\pi^{-1}(\tilde{\gamma})}\Phi(e_{1}(\tau))\tilde{\omega}=\int_{[\gamma]}\pi_{1\ast}\Phi(e_{1}(\tau))=\int_{[\gamma]}\Psi\tau.$

### v2

@@ 3. Action on holomorphic differentials
We have seen that by passing through a specific type of $2$-currents instead of directly using the pullback $\phi^{*}$, the linear map $\Phi:\mathcal{D}^{\prime}_{\text{tr}}(SM)\to\mathcal{D}^{\prime}_{\text{tr}}(\widetilde{S}M)$ defined through the diagram (2.3) sends distributional solutions to the transport equation of one thermostat to those of the second. In this section, we want to show that $\Phi$ can also be seen as acting on holomorphic differentials from one complex surface to another when $\lambda$ is of Fourier degree $\leq 1$ and the attenuated tensor tomography problem of order $1$ is satisfied.
3.1. Action on fibrewise holomorphic distributions
We start by studying the action of $\Phi$ on the subspace $\mathcal{D}^{\prime}_{\text{tr},+}(SM)$ defined in equation (2.6). This will require some microlocal analysis.
We introduce $\mathcal{C}\subset\left\{(v,\xi)\in T^{*}(SM)\mid\xi(F(v))=0\right\}$, the closed cone enclosed by $E_{s}^{\ast}$ and $E_{u}^{\ast}$ in the half-space $\left\{(v,\xi)\in T^{*}(SM)\mid\xi(V(v))\geq 0\right\}$. See Figure 1.
@@ 3.1. Action on fibrewise holomorphic distributions
Let $(M,g,\lambda)$ be an Anosov thermostat. If $u\in\mathcal{D}^{\prime}_{\textup{tr},+}(SM)$, then we have $\textup{WF}(u)\subset\mathcal{C}$ and $u_{k}\in\Omega_{k}$ for all $k\in\mathbb{Z}$.
@@
The argument is essentially the same as that of [GLP25, Lemma 2.5]. Let us give the details.
@@
$\text{WF}(u)=\text{WF}(S_{+}u)\subset\left\{(v,\xi)\in T^{*}(SM)\mid\xi(V(v))\geq 0\right\}.$
Given that $(F+V\lambda)u=0$, elliptic regularity tells us that
$\text{WF}(u)\subset\Sigma.$
By propagation of singularities for real principal type differential operators (see [Hör09, Theorem 26.1.1]), we further know that $\text{WF}(u)$ is invariant by the symplectic lift to $T^{*}(SM)$ of the flow $\{\varphi_{t}\}_{t\in\mathbb{R}}$. Given the Anosov property, the maximal flow-invariant subset of $T^{*}(SM)$ contained in $\Sigma\cap\left\{(v,\xi)\in T^{*}(SM)\mid\xi(V(v))\geq 0\right\}$ is precisely $\mathcal{C}$, so this gives us the first claim.
For the second claim, recall that $u_{k}=(2\pi)^{-1}\pi_{k}^{\ast}\pi_{k\ast}u$. The pushforward operator $\pi_{k\ast}$ only selects the wavefront set in $(\mathbb{R}X)^{\ast}\oplus\mathbb{H}^{*}$ (see [FT99, Proposition 11.3.3]), which is empty given that $\mathcal{C}\cap((\mathbb{R}X)^{\ast}\oplus\mathbb{H}^{*})=\{0\}$ by property (2.11). Therefore, $u_{k}\in\Omega_{k}$. ∎
We will also need the following lemma with the same proof as [GLP25, Lemma 3.3].
@@
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$ be Anosov thermostats. Suppose that there exists a smooth orbit equivalence $\phi:\widetilde{S}M\to SM$ isotopic to the identity between them. Then, $\phi$ preserves the natural orientation of the weak unstable subbundle $\mathbb{R}F\oplus E_{u}$, namely that given by the global frame $(F,Y^{u})$, where $Y^{u}$ is defined in equation (2.10).
Armed with this, we can show that $\Phi$ maps fibrewise holomorphic distributional solutions to the transport equation of one thermostat to those of the second. For this step of the proof, however, we restrict to thermostats where $\lambda$ has Fourier degree $\leq 1$ and the attenuated tensor tomography problem of order $1$ is satisfied.
@@
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$, with $\lambda$ and $\tilde{\lambda}$ of Fourier degree $\leq 1$, be Anosov thermostats satisfying the attenuated tensor tomography problem of order $1$. Suppose there exists a smooth orbit equivalence $\phi:\widetilde{S}M\to SM$ isotopic to the identity between them. Then, the map $\Phi$ defined through the diagram (2.3) yields a $\mathbb{C}$-linear isomorphism
$\Phi:\mathcal{D}^{\prime}_{\textup{tr},+}(SM)\to\mathcal{D}^{\prime}_{\textup{tr},+}(\widetilde{S}M).$
@@
Since $d\phi^{\top}$ maps connected sets to connected sets, $E_{u}^{*}$ to $\widetilde{E}_{u}^{*}$, and $E_{s}^{*}$ to $\widetilde{E}_{s}^{*}$, $d\phi^{\top}(\mathcal{C})$ must be one of the four cones depicted on the right of Figure 1 (inside the characteristic set $\widetilde{\Sigma}$). It follows that $d\phi^{\top}(\mathcal{C})=\pm\widetilde{\mathcal{C}}$ because any other cone would entail that $\phi$ reverses the orientation, which is impossible since it is assumed to be isotopic to the identity. If $d\phi^{\top}(\mathcal{C})=-\widetilde{\mathcal{C}}$, then $\phi$ would flip the orientation of the weak unstable leaves, which contradicts Lemma 3.2. Therefore, $d\phi^{\top}(\mathcal{C})=\widetilde{\mathcal{C}}$.
Let $u\in\mathcal{D}^{\prime}_{\textup{tr},+}(SM)$ and $\tilde{u}\coloneqq\Phi u$. By Lemma 3.1, we know that $\text{WF}(u)\subset\mathcal{C}$. By Lemma 2.4, we thus know that $\text{WF}(\tilde{u})=d\phi^{\top}(\text{WF}(u))\subset\widetilde{\mathcal{C}}$. Then, $\widetilde{S}_{-}\tilde{u}\in\mathcal{C}^{\infty}(\widetilde{S}M)$ and, since $(\widetilde{F}+\widetilde{V}\tilde{\lambda})\tilde{u}=0$, we also have that
$(\widetilde{F}+\widetilde{V}\tilde{\lambda})\widetilde{S}_{-}\tilde{u}$ $=[\widetilde{F}+\widetilde{V}\tilde{\lambda},\widetilde{S}_{-}]\tilde{u}$
$=(\tilde{\eta}_{+}+\tilde{\lambda}_{1}\widetilde{V}+i\tilde{\lambda}_{1})\tilde{u}_{0}-(\tilde{\eta}_{-}+\tilde{\lambda}_{-1}\widetilde{V}-i\tilde{\lambda}_{-1})\tilde{u}_{1}.$
Since $(M,\tilde{g},\tilde{\lambda})$ satisfies the tensor tomography problem of order $1$ by assumption, it follows that $\widetilde{S}_{-}\tilde{u}$ is of Fourier degree 00. Hence, $\tilde{u}$ is fibrewise holomorphic. The fact that $\Phi$ is an isomorphism is then clear as it admits an inverse, namely the map associated to $(\phi^{-1})^{*}$ by the diagram (2.3). ∎
@@ 3.2. Extension operator
is well-defined. Indeed, if $u\in\mathcal{D}^{\prime}_{\text{tr},+}(SM)$, then $Xu+V(\lambda u)=(F+V\lambda)u=0$, which means that $(Xu)_{0}=0$ and hence $\eta_{-}u_{1}=0$. By Lemma 2.5, this is equivalent to ¯$\overline{\partial}\pi_{1\ast}u=0$.
@@
$\Psi:H^{0}_{J}(M,\kappa)\to H^{0}_{\widetilde{J}}(M,\kappa)$
by the following commutative diagram.
${\mathcal{D}^{\prime}_{\text{tr},+}(SM)}$${\mathcal{D}^{\prime}_{\text{tr},+}(\widetilde{S}M)}$${H^{0}_{J}(M,\kappa)}$${H^{0}_{\widetilde{J}}(M,\kappa)}$$\scriptstyle{\Phi}$$\scriptstyle{\pi_{1\ast}}$$\scriptstyle{e_{1}}$$\scriptstyle{\Psi}$ (3.2)
@@ 3.3. Period preservation
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$, with $\lambda$ and $\tilde{\lambda}$ of Fourier degree $\leq 1$, be Anosov thermostats satisfying the attenuated tensor tomography problem of order $1$. Then, the $\mathbb{C}$-linear map
$\Psi:H^{0}_{J}(M,\kappa)\to H^{0}_{\widetilde{J}}(M,\kappa)$
@@
Recall that there is a push-forward map $\pi_{\ast}:\mathcal{C}^{\infty}(SM,\Omega^{2}(SM))\to\mathcal{C}^{\infty}(M,\Omega^{1}(M))$ given by integration along fibres. It satisfies $d\pi_{\ast}=\pi_{\ast}d$ (see [BT82, Proposition 6.14]), and it extends to currents. By [BT82, Proposition 6.15], we have the projection formula
@@
Let $(M,g,\lambda)$ be a thermostat and let $\omega$ be the $2$-form on $SM$ defined in equation (2.2). For any $u\in\mathcal{D}^{\prime}(SM)$, we have
@@
It suffices to establish the claim for $u\in\mathcal{C}^{\infty}(SM)$. Recall that $\omega\coloneqq\iota_{F}\mu$. A quick computation then yields
@@
$\pi_{\ast}(u\omega)_{x}(w)=\int_{S_{x}M}\iota_{\tilde{w}}(u\omega),$
where $\tilde{w}\in T(SM)$ is a lift of $w$ under $d\pi$. We take $\tilde{w}=(w,0)$, that is, no component in the vertical subbundle $\mathbb{V}=\ker d\pi$. Then, since $\psi(\tilde{w})=0$, we get
$\pi_{\ast}(u\omega)_{x}(w)=\int_{S_{x}M}\iota_{\tilde{w}}(\beta)u\psi.$
@@
where in the last equality, we used equation (2.4). In terms of $1$-forms, since we have $u_{k}=(2\pi)^{-1}\pi_{k}^{*}\pi_{k\ast}u$, we proved that
@@
We conclude by applying $\star$ to both sides. ∎
@@
Let $(M,g,\lambda)$ be an Anosov thermostat and let $\gamma$ be a closed thermostat geodesic. For any $u\in\mathcal{D}^{\prime}_{\textup{tr}}(SM)$, the pairing $\langle\pi^{-1}(\gamma),u\omega\rangle$ is well defined and
@@
By the wavefront set calculus, the pairing $\langle\pi^{-1}(\gamma),u\omega\rangle$ is well defined whenever
@@
(see [Hör03, Corollary 8.2.7] for instance). Since $\mathbb{V}\subset T(\pi^{-1}(\gamma))$, the conormal bundle $N^{*}(\pi^{-1}(\gamma))$ consists of a line contained in $(\mathbb{R}X)^{*}\oplus\mathbb{H}^{*}$. Lemma 3.1 and property (2.11) then tell us that the intersection with $\text{WF}(u)$ is indeed empty. It follows that the pairing $\langle\pi^{-1}(\gamma),u\omega\rangle$ is well defined and extends the pairing computed for $u\in\mathcal{C}^{\infty}(SM)$.
We can then apply the projection formula (3.3) and Lemma 3.5. By Lemma 2.3, the $2$-current $u\omega$ is closed if $u\in\mathcal{D}^{\prime}_{\text{tr}}(SM)$, so $\star(\pi_{-1\ast}u+\pi_{1\ast}u)$ is also closed given that $\pi_{*}d=d\pi_{*}$, which implies that its integral only depends on the homology class $[\gamma]$. ∎
As $\ell_{1}\star=-V\ell_{1}$, for any $u\in\mathcal{D}^{\prime}(SM)$, we may write
@@
Therefore, if $\tau\in H^{0}_{J}(M,\kappa)$, $[\gamma]\in H_{1}(M;\mathbb{Z})$ and $\gamma$ is any thermostat geodesic whose homology class is $[\gamma]$, Lemma 3.6 gives us
@@
Let $[\gamma]\in H_{1}(M;\mathbb{Z})$ and let $\gamma$, $\tilde{\gamma}$ be two thermostat geodesics with respect to $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$, respectively, whose homology class is $[\gamma]$. Since $\phi$ is isotopic to the identity, we know that $[\phi(\pi^{-1}(\tilde{\gamma}))]=[\pi^{-1}(\gamma)]$ in $H_{2}(SM;\mathbb{Z})$.
We claim that the pairing $\langle\phi(\pi^{-1}(\tilde{\gamma})),e_{1}(\tau)\omega\rangle$ is well defined. In light of condition (3.5) and Lemma 3.1, which tell us that $\text{WF}(e_{1}\tau)\subset\mathcal{C}$, it is enough to show that
$N^{*}(\phi(\pi^{-1}(\tilde{\gamma})))\cap\mathcal{C}=\{0\}.$ (3.7)
We have seen in the proof of Proposition 3.3 that $d\phi^{\top}(\mathcal{C})=\widetilde{\mathcal{C}}$, so combining this observation with the property
$N^{*}(\phi(\pi^{-1}(\tilde{\gamma})))=(d\phi^{\top})^{-1}(N^{*}(\pi^{-1}(\tilde{\gamma})))$
allows us to rewrite condition (3.7) as
$N^{*}(\pi^{-1}(\tilde{\gamma}))\cap\widetilde{\mathcal{C}}=\{0\}.$
Again, since $\widetilde{\mathbb{V}}\subset T(\widetilde{S}M)$, we know that the conormal bundle $N^{*}(\pi^{-1}(\tilde{\gamma}))$ consists of a line in $(\mathbb{R}\widetilde{X})^{*}\oplus\widetilde{\mathbb{H}}^{*}$, so it trivially intersects the closed cone $\widetilde{\mathcal{C}}$, as desired.
The $2$-current $e_{1}(\tau)\omega$ is closed by Lemma 2.3. By the Hodge decomposition theorem, we may hence write $e_{1}(\tau)\omega=\sigma+df$ for some harmonic $2$-current $\sigma$ and $1$-current $f$ with $\text{WF}(f)=\text{WF}(e_{1}\tau)$. Thanks to the wavefront set condition, the same argument as for $e_{1}(\tau)\omega$ then shows that both pairings $\langle\pi^{-1}(\gamma),df\rangle$ and $\langle\phi(\pi^{-1}(\tilde{\gamma})),df\rangle$ are well defined. They must be equal to 00 since $df$ is exact. We thus get
@@
where in the second equality, we have used the facts that $\sigma$ is harmonic and that $[\pi^{-1}(\gamma)]=[\phi(\pi^{-1}(\tilde{\gamma}))]$ in $H_{2}(SM;\mathbb{Z})$.
We can now use equation (3.6) and unravel the definitions to obtain
@@
$=2i\int_{\pi^{-1}(\tilde{\gamma})}\Phi(e_{1}\tau)\tilde{\omega}=\int_{[\gamma]}\pi_{1\ast}\Phi(e_{1}\tau)=\int_{[\gamma]}\Psi\tau.\qed$